OpenStudy (infinexa):

A number is written with the following factorization 2 x 3 x 5^4 x 8 x 11^2 . Is this factorization a prime factorization? Explain why or why not. If it is not correct, give the correct prime factorization of the number.

OpenStudy (infinexa):

@Directrix @Nnesha @WanderingGirl

OpenStudy (rizags):

Do you know what a prime number is?

OpenStudy (infinexa):

yes

OpenStudy (rizags):

are there any numbers in your list of factors that are not prime?

OpenStudy (rizags):

(not including exponents of course)

OpenStudy (infinexa):

what list of factors

OpenStudy (rizags):

Up there in your question, you have a "list" of factors. I will state them explicitly: 2, 3, 5, 5, 5, 5, 8, 11, 11. Which, if any, are not prime?

OpenStudy (infinexa):

8 is not prime

OpenStudy (rizags):

correct. What is the prime factorization of 8? do you know?

OpenStudy (infinexa):

2^3

OpenStudy (rizags):

good. Now we can rewrite your list of factors as: 2, 2, 2, 2, 3, 5, 5, 5, 5, 11, 11. Do you understand how i did this?

OpenStudy (infinexa):

mmm not really

OpenStudy (rizags):

well, since 8 = 2^3, we can say that 8=2*2*2, so instead of an 8 in your list of prime factors, I instead inserted three twos. Make sense?

OpenStudy (infinexa):

oh yes i understand

OpenStudy (rizags):

ok, now all that is left is to multiply and condense the factors. do you know how to do that?

OpenStudy (infinexa):

i can multiply the factors but what is condense the factors

OpenStudy (rizags):

condense is to write them in exponent form. DON'T actually multiply them, just write them in exponent form. for example, {2, 2, 2, 2} is just 2^4

OpenStudy (infinexa):

yes

OpenStudy (rizags):

can you do that?

OpenStudy (infinexa):

yes

OpenStudy (rizags):

ok tell me what you get

OpenStudy (infinexa):

2^4 x 3 x 5^4 x 11^2

OpenStudy (rizags):

\[\huge 2^4\times3\times5^4\times11^2\]

OpenStudy (rizags):

This is correct

OpenStudy (infinexa):

yes

OpenStudy (infinexa):

so it is a prime factorization

OpenStudy (rizags):

Yes this is the prime factorization

OpenStudy (infinexa):

thank you

OpenStudy (infinexa):

it says explain

OpenStudy (infinexa):

can you explain why

OpenStudy (rizags):

No problem. Medal Please! Your explanation is that all of the factors listed are primes raised to exponents, thus the product represents the prime factorization

OpenStudy (infinexa):

thank you and here is a medal

OpenStudy (rizags):

great thanks!